Rosa is 18 years old and Arthur is 24 years old. How many years ago was Arthur three times as old as Rosa?
Answer by ankor@dixie-net.com(12695)

osa is 18 years old and Arthur is 24 years old. How many years ago was Arthur three times as old as Rosa?
:
Let x = number of years ago that this statement was true
:
A's age x yrs ago = 3 times R's age x yrs ago
:
24 - x = 3(18 - x)
:
24 - x = 54 - 3x; multiplied what's inside the brackets
:
-x + 3x = 54 - 24
:
2x = 30
:
x = 15 yrs ago. You can check and see if this is true: